The mgf of X is Mx(t) = (¿e-* + ¿e Find P(X > -3|X < 2). -3t 0.86 0.17 0.91 none of the answers provided here 0.29 2) 5/6
Answer #1

expanding above :M(t) =(1/36)*e-6t +(25/36)e4t +(10/36)e-t  

therefore

P(X=-6)=1/36

P(X=-1) =10/36

P(X=4)=25/36

P(X>-3|X<2) =P(-3<X<2)/P(X<2) =(10/36)/(11/36)=10/11 =0.91